OpenStudy (anonymous):

Simplify the radical expression. √72

OpenStudy (anonymous):

a. \[6\sqrt{2}\] b. 2 c. 12

OpenStudy (anonymous):

\[\sqrt{72} = \sqrt{6\times6\times2} = \sqrt{6^{2}\times2} = 6\sqrt{2}\]

OpenStudy (anonymous):

o thanks...

OpenStudy (anonymous):

anytime